OpenStudy (anonymous):

properties of rational exponents: (a^2)^3

OpenStudy (anonymous):

\[(a ^{2})^{3}\]

OpenStudy (anonymous):

write out what it means... then you could probably figure out the shortcut if you do enough of these: (a*a)^3 = (a*a) * (a*a) * (a*a) = a*a*a*a*a*a =

OpenStudy (anonymous):

so it would be \[a ^{6}\]

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

do you see a shortcut way of doing this?

OpenStudy (anonymous):

multiplying the powers together?

OpenStudy (anonymous):

exactly... :) now you don't have to do the long way..

OpenStudy (anonymous):

awesome, Thank you. :)

OpenStudy (anonymous):

yw